OpenStudy (anonymous):

f(x)=x^3+8 find all zeros by factoring each function.

OpenStudy (anonymous):

hi

jhonyy9 (jhonyy9):

x=2

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

-2

jhonyy9 (jhonyy9):

is right

OpenStudy (anonymous):

they has 3 different result, yours is one (-2)

jhonyy9 (jhonyy9):

3 result ???

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

looke x^3 +8=0 x^3 =-8 = -2^3

OpenStudy (anonymous):

{-2, 1+i sqrt 3, 1-i sqrt3

OpenStudy (anonymous):

I think 3^+8= (x+2)(x^2-2x+4)

OpenStudy (anonymous):

so x=-2 and x^2-2x+4 will give other two results

OpenStudy (anonymous):

use squardric formular to do it

OpenStudy (anonymous):

I think you gone

jhonyy9 (jhonyy9):

yes is right and you need to calculate x from x+2 from this will get x=-2 and from x^2 -2x +4 2+/- radical (4-16) 2(1+/- i radical 3) x_1,2 = ------------------ = ----------------- = 1+/- i radical 3 2 2 radical hope that you know equal sqrt

jhonyy9 (jhonyy9):

hope that this is clearly

OpenStudy (anonymous):

yesbu i in this case is 1^2?

OpenStudy (anonymous):

i in this case is (-1)^2?

jhonyy9 (jhonyy9):

sorry what is this what you have wrote ?

OpenStudy (anonymous):

I don't understand why you use i-radial 3. What is i ?

OpenStudy (anonymous):

I think i = (-1)^2

jhonyy9 (jhonyy9):

if you see there is x_1,_2 = (2+/- sqrt (4-16))/2 and here from sqrt(-12) you get 2i sqrt 3 but after you simplified with 2 will remain this resulte

jhonyy9 (jhonyy9):

now is clearly ?

OpenStudy (anonymous):

so sqrt (-2)^* 3 = i sqrt 3

OpenStudy (anonymous):

yes it is clear

jhonyy9 (jhonyy9):

sqrt(-2)^3 = sqrt (-2^2)*2 =2i sqrt 2 because i^2 = -1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i understand now

jhonyy9 (jhonyy9):

ok bye

OpenStudy (anonymous):

thanks for ur help

jhonyy9 (jhonyy9):

was my pleasure